OpenStudy (anonymous):

a man left 1/2 of his money to his wife, 1/6 to his daughter, and remainder on amount of 15,000 to his son. how much money did he leave?

OpenStudy (anonymous):

assume it to be x

OpenStudy (anonymous):

the money he leave?

OpenStudy (anonymous):

total money, the amount u need to find

OpenStudy (anonymous):

how will u get the value of 1/2 and 1/6 (the fraction part) ???

OpenStudy (anonymous):

x/2 and x/6

OpenStudy (anonymous):

k. tnx

OpenStudy (anonymous):

a bit confused........ ........but........u help a lot

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{6}=\frac{2}{3}\] right? so the remaining amount is one third of the money. you are told that one third of the money is 15,000 so you can figure out what the total amount is

OpenStudy (callisto):

It's the third time I saw this question today! http://openstudy.com/study#/updates/4ff42169e4b01c7be8c78d4f

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

>he left 12500 to his wife. >7500 to his daughter >so all together he left 30000